what atomic orbitals are hybridized for

PO4^3-

ClO3^-

ClO4^-

NO2^-

NO3^-

CO2

SF4^2-

SO4^2-

SO3^2-

AsO4^3

SnCl2

CH4

NH3

H2O

PCl5

SF4

BrF3

XeF2

SF6

IF5

XeF4

BF3
Answer #1

We need to analyze to how many groups the central atom is bonded and lone pairs, with that number, we can predict is hybridization:

PO4^3- = sp3

ClO3^- = sp3

ClO4^- = sp3

NO2^- = sp2

NO3^- = sp2

CO2 = sp

SF4^2- = sp3d2

SO4^2- = sp3

SO3^2- = sp3

AsO4^3 = sp3

SnCl2 = sp2

CH4 = sp3

NH3 = sp3

H2O = sp3

PCl5 = sp3d

SF4 = sp3d

BrF3 = sp3d

XeF2 = sp3d

SF6 = sp3d2

IF5 = sp3d2

XeF4

BF3